OpenStudy (anonymous):

Write the equation of the line with a slope of 3/7 passing through the point (9, -2)

OpenStudy (anonymous):

Use the point-slope form: \[y-y_{0}=m(x-x _{0})\] where m = 3/7, x0=9 and y0=-2 y+2=3/7(x-9) 7(y+2)=3x-27 7y+14=3x-27 In standard form: 3x-7y-41=0

OpenStudy (anonymous):

There are two basic ways to write a line: If have you a slope and a y-intercept: y=mx+b If you have 2 points though: (y-k)=m(x-h) That second one can look longer, but if you remember that x is before y in the alphabet and h is before k, it is actually really helpful. So: (y-k)=m(x-h) y and x stay there, you just substitute in k, h, and m. m is your slope: 3/7 k and h are (h,k) you point: (0,-2) Therefore: (y-(-2))=(3/7)(x-0) y+2=(3/7)x y+2=3x/7 y=(3x/7)-2